• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen/gilman

    Log entry time 14:24:07 on April30,2007

    Entry number 200635

    keyword=cryotarget (He4) cool-down completed

    After CHL stabalized, MCC informed that we could cool the target.
    Closed the vent valve. Started cool-down. Switched to a new bottle of He4 gas when teh current bottle became empty.
    The cool-down was completed. The current oprating condition is
    8.5K, 201 psi, 60 Hz. With 15.5g/s we have 340watts. Will need to
    increase coolant flow before ready for beam.


    FIGURE 1